1. Use the commutative property to write an expression that is equivalent to 3 +(-5)

2. show whether -7+2 is the same as 7+(-2)

Answer by jim_thompson5910(35256)

# 1
3+(-5) = -5 + 3

-------------------------------------------------------

# 2
-7 + 2 does NOT equal 7 + (-2) since -7+2 = -5 and 7 + (-2) = 5